LEAST SQUARES FITTING OF PERTURBED VIBRATIONAL POLYADS NEAR THE ISOMERIZATION BARRIER IN THE S_{1} STATE OF C_{2}H_{2}

\underline{A. J. MERER}, Department of Chemistry, University of British Columbia, Vancouver, B.C., Canada V6T 1Z1 AND Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.; J. H. BARABAN, P. B. CHANGALA, and R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139, USA.

The S_{1} electronic state of acetylene has recently been shown to have two potential minima, corresponding to cis- and trans-bent structures. The trans-bent isomer is the more stable, with the cis-bent isomer lying about 2670 cm^{−1} higher; the barrier to isomerization lies roughly 5000 cm^{−1} above the trans zero-point level. The “isomerization coordinate” (along which the molecule moves to get from the trans minimum to the barrier) is a combination of the ν_{3} (trans bending) and ν_{6} (cis bending) vibrational normal coordinates, but the spectrum is very confused because the ν_{6} vibration interacts strongly with the ν_{4} (torsion) vibration through Coriolis and Darling-Dennison resonances. Since the ν_{4} and ν_{6} fundamental frequencies are almost equal,^{a} the bending vibrational structure consists of polyads. At low vibrational energies the polyads where these three vibrations are excited can be fitted by least squares almost to experimental accuracy with a simple model of Coriolis and Darling-Dennison interactions, but at higher energies the huge x_{36} cross-anharmonicity, which is a symptom that the levels are approaching the isomerization barrier, progressively destroys the polyad structure; in addition the levels show an increasing even-odd staggering of their K-rotational structures, as predicted by group theory. It is not possible to fit the levels near the barrier with a simple model, though some success has been achieved with extended models. Progress with the fitting of the polyads near the barrier will be reviewed.

aA. L. Utz, J. D. Tobiason, E. Carrasquillo M., L. J. Sanders and F. F. Crim, J. Chem. Phys. \underline{98}, 2742, 1993.